What is the kinetic energy of the continent?


1
Expert's answer
2021-11-19T10:51:24-0500

The mass of the North American continent, for instance, is estimated to be "2.35095\cdot10^{23}"  kg. It moves at the rate of 3.8 cm/year. The kinetic energy is


"K=0.5(2.35095\\cdot10^{23})\\left(\\frac{3.8}{100\u22c5365\u22c53600\n\n\u200b\n }\\right)^2=1706.74 \\ J"